OpenStudy (anonymous):

the polynomial x^5-3x^4+2x^3-2x^2+3x+1 is denoted by f(x). when f(x) is divided by (x^2+1), the remainder is 2x. find all the real roots of the equation f(x)=2x

OpenStudy (anonymous):

\[f(x)=(x^2+1)(something)+2x\]

OpenStudy (anonymous):

\(x^2+1\) has no real roots, so it will have to be the real roots of the "something"

OpenStudy (anonymous):

x^3-3x^2+x+1=something?

OpenStudy (anonymous):

i have no idea, but i can check

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

great however you got it, it is right

OpenStudy (rajee_sam):

they are saying find the real roots of the equation f(x) = 2x. What does that mean?

OpenStudy (anonymous):

so your job is now to find the roots of \(x^3-3x^2+x+1\)

OpenStudy (anonymous):

i also dont know

OpenStudy (anonymous):

which is not that hard, because one of the roots is 1

OpenStudy (anonymous):

so you can factor as \((x-1)(x^2-2x-1)\) and solve for the second part via the quadratic formula

OpenStudy (anonymous):

what is the meaning of f(x)=2x?

OpenStudy (anonymous):

\[f(x)=2x\] means exactly that we know one answer is 1, this means \(f(1)=2\)

OpenStudy (anonymous):

find a number for which the function evaluated at the number is the same as two times the number

OpenStudy (anonymous):

for example if \(f(3)=6\) what would be a solution it is not, that is just an example

OpenStudy (rajee_sam):

f(x) is not ( x³ - 3x² + x +1 ) it is x^5 - 3x^4 + 2x³ -2x² +3x +1

OpenStudy (anonymous):

right, and \(f(1)=1-3+2-2+3+1=2\) and also \(2\times 1=2\)

OpenStudy (anonymous):

which means if \(x=1\) then \(f(x)=2x\)

OpenStudy (rajee_sam):

I understand what you are doing here but just a bit confused with the wording of the problem that is all

OpenStudy (anonymous):

you still have to find the roots of \(x^2-2x-1\)

OpenStudy (rajee_sam):

sort of get it now

OpenStudy (anonymous):

yeah it is kind of confusing you can think of it as \(f(x)=x^5 - 3x^4 + 2x³ -2x² +3x +1 \) and \(g(x)=2x\) solve \[f(x)=g(x)\]

OpenStudy (anonymous):

or for that matter solve \[x^5 - 3x^4 + 2x³ -2x² +3x +1 =2x\]

OpenStudy (rajee_sam):

that is exactly what I did and we get x^5 - 3x^4 + 2x^3 - 2x^2 + x +1 = 0